OpenStudy (anonymous):

r= 2/3t+v , solve for t Show work please, I'm stressing out on some of these problems... thanks :)

OpenStudy (anonymous):

r=2/3t+v so to solve for T.. first take minus the V on both sides like this v-r=2/3t you following this??

OpenStudy (anonymous):

next because we are solving for T we divide by 2/3 to get T my itself. so we have v-r/2/3=t

OpenStudy (anonymous):

hope this helps...

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

not a problem any questions before I leave?

OpenStudy (anonymous):

I have one last problem thats bugging me and that is -14n+q=rt-4n, solve for n.

OpenStudy (anonymous):

first what do you know already

OpenStudy (anonymous):

do I need to subtract rt from both sides?

OpenStudy (anonymous):

yes you can do that, but we to collect a like terms first

OpenStudy (anonymous):

Oh! The -14n and the -4n can go together.

OpenStudy (anonymous):

yes, but be careful here what happens with the sign when it crosses the equal? recall + becomes - and - becomes +

OpenStudy (anonymous):

So, 18n+g=rt ?

OpenStudy (anonymous):

that is correct

OpenStudy (anonymous):

so now what do we need to get N standing alone?

OpenStudy (anonymous):

hint look at the g.

OpenStudy (anonymous):

subtract the g on both sides, leaving me with 18n=rt-g

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then divide by 18 to get n=rt-g/18 ?

OpenStudy (anonymous):

Nice job

OpenStudy (anonymous):

Thank you SO much! I understand it now! :D You're a lifesaver.

OpenStudy (anonymous):

I want to be sure you understand so ask questions if you have any?

OpenStudy (anonymous):

Okay, I have another problem that I know how to start off but not finish. ax+z=aw -y +y +y And that gives me ax+z+y=aw And thats where I'm not sure where to go from there...

OpenStudy (anonymous):

what are we solving for or what is the problem asking for?

OpenStudy (anonymous):

whoops. the equation is ax+z=aw-y, solve for a. You add y on both sides giving you ax+z+y=aw. And then I don't know what else to do.

OpenStudy (anonymous):

okay this one is more tricky, but do able

OpenStudy (anonymous):

first let's look at the a terms here? are alike? no, but what we can do is them on side.

OpenStudy (anonymous):

so you can do this ax+z+y=aw take the ax-aw and than take -z-y so you should have the following ax-aw=-z-y

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

this is where it is tricky what can we do with the a's here?

OpenStudy (anonymous):

add the 2 a's to get 2a(wx)=-z-y ?

OpenStudy (anonymous):

try factoring an a.

OpenStudy (anonymous):

did you factor an A?

OpenStudy (anonymous):

I don't understand again..

OpenStudy (anonymous):

I have it 2a(wx)=-z-y ? from this point take the wx and divide by -z-y so therefore you would 2a=-z-y/wx next divide by 2 to get the a alone so you should have A=-z-y/2wx.

OpenStudy (anonymous):

Okay thank you so much! I understand it now! :D

OpenStudy (anonymous):

really make sure you understand ask questions plz if something isn't clear?